# Refining local-type primordial non-Gaussianity: Sharpened $b_{\phi}$ constraints through bias expansion

### v1

@@ III.2.1 Constructing fair galaxy samples
We have already learned that the halo properties that have a strong dependence on the value of $b_{\phi}$ are the halo mass and assembly bias parameters such as concentration and environment. For this reason, in designing the galaxy samples for this study, we adopted the best-fit parameters of the extended HOD analysis of Ref. (2024MNRAS.530..947Y, ), exploring modifications to six of the model parameters, namely: $\log M_{\rm cut}$, $\log M_{1}$, the two main parameters controling the halo masses entering the sample, $A_{\rm c}$, $A_{\rm s}$, the concentration assembly bias parameters, and $B_{\rm c}$, $B_{\rm s}$, the environment assembly bias parameters. Specifically, we vary the mass parameters within 0.2 dex of their best-fit values, the concentration ones between -2 and 2, and the environment ones between -0.5 and 0.5. We note that some of these variations appear to be ruled out by the tight constraints placed on the data, but at the same time, assembly bias ornamentations are currently not too well constrained.
@@ III.2.3 Best-fit parameters
The $\beta$ coefficient values from this table along with the measured bias parameters $b_{X}$ can be used jointly to predict $b_{\phi}$ and serve as prior for a PNG analysis. It is reassuring that the bestfit $\beta$ coefficients for the two LRG samples are very similar, whereas the quasar sample yields slightly different values for $\beta$. We note that these values are specific to analysis perform with the AbacusSummit suite and AbacusHOD prescription. In addition, when folding in our linear regression model into a PNG fit to real data, one would also need to marginalize over the error on these coefficients. Therefore, our recommendation is to perform the $b_{\phi}$ predictions on the fly with error bars when using AbacusSummit and otherwise, conduct a thorough recalibration of the fits.
@@ Acknowledgements.
BH is supported by the Miller Institute for Basic Science. SF is supported by Lawrence Berkeley National Laboratory and the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Additional computations in this work were performed at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation.

### v2

@@ III.2.1 Constructing fair galaxy samples
We have already learned that the halo properties that have a strong dependence on the value of $b_{\phi}$ are the halo mass and assembly bias parameters such as concentration and environment. For this reason, in designing the galaxy samples for this study, we adopted the best-fit parameters of the extended HOD analysis of Ref. (2024MNRAS.530..947Y, ), exploring modifications to six of the model parameters, namely: $\log M_{\rm cut}$, $\log M_{1}$, the two main parameters controlling the halo masses entering the sample, $A_{\rm c}$, $A_{\rm s}$, the concentration assembly bias parameters, and $B_{\rm c}$, $B_{\rm s}$, the environment assembly bias parameters. Specifically, we vary the mass parameters within 0.2 dex of their best-fit values, the concentration ones between -2 and 2, and the environment ones between -0.5 and 0.5. We note that some of these variations appear to be ruled out by the tight constraints placed on the data, but at the same time, assembly bias ornamentations are currently not too well constrained.
@@ III.2.3 Best-fit parameters
The $\beta$ coefficient values from this table along with the measured bias parameters $b_{X}$ can be used jointly to predict $b_{\phi}$ and serve as prior for a PNG analysis. It is reassuring that the best-fit $\beta$ coefficients for the two LRG samples are very similar, whereas the quasar sample yields slightly different values for $\beta$. We note that these values are specific to the analysis performed with the AbacusSummit suite and AbacusHOD prescription. In addition, when folding in our linear regression model into a PNG fit to real data, one would also need to marginalize over the error on these coefficients. Therefore, our recommendation is to perform the $b_{\phi}$ predictions on the fly with error bars when using AbacusSummit and otherwise, conduct a thorough recalibration of the fits.
@@ Acknowledgements.
We thank Stephen Chen, Nick Kokron, Elisabeth Krause, Uroš Seljak, James Sullivan, and Martin White for very helpful discussions and comments on the draft. BH is supported by the Miller Institute for Basic Science. SF is supported by Lawrence Berkeley National Laboratory and the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Additional computations in this work were performed at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation.